OpenStudy (ajanijones):

HELPEEE. A cylinder has a base with a radius of 7 cm and a height of 16 cm. What is the volume of the cylinder? (Use π = 3.14.) A. 703.36 cm3 B. 351.68 cm3 C. 2461.76 cm3 D. 615.44cm3

OpenStudy (anonymous):

Vcylinder = π.r^2.h Vcylinder = 3,14(7^2).16 Vcylinder = 2461,76

OpenStudy (ajanijones):

So my answer's C right.

OpenStudy (anonymous):

c